OpenStudy (anonymous):

the function f(x)=4/x+2 is one to one find an equation for f^-1(x) the inverse function f^-1(x) please help

OpenStudy (anonymous):

You just need to use the definition of inverse function. A function \[f^{-1}(x)\]is an inverse to the function \[f(x)\]if\[f(f^{-1}(x))=f^{-1}(f(x))=x\]So this is what you do. For \[f(x)=\frac{4}{x}+2\]applying the definition,\[f(f^{-1}(x))=\frac{4}{f^{-1}(x)}+2=x\]So\[\frac{4}{f^{-1}(x)}=x-2\rightarrow f^{-1}(x)=\frac{4}{x-2}\]

OpenStudy (anonymous):

ima guess this is \[f(x)=\frac{4}{x+2}\]

myininaya (myininaya):

\[y=\frac{4}{x+2}\] \[y(x+2)=4\] \[xy+2y=4\] \[xy=4-2y\] \[x=\frac{4-2y}{y}\] \[f^{-1}(x)=\frac{4}{x}-2\]

OpenStudy (anonymous):

myininaya help me here will you please?